Find the value of X & Y in the given linear equations 3x+2y=14 and 5x+4y =26 .


Eco255 test questions
1
Expert's answer
2020-03-26T10:43:55-0400

"3x+2y=14 | \u00d72"

"6x + 4y = 28"

-

"5x+4y = 26"

=

"x = 2,"

"3\u00d72 + 2y = 14,"

"2y = 8,"

"y = 4."